If z^2=x^2+y^2 with z>0, dx/dt=2, and dy/dt=5, find dz/dt when x=5 and y=12.

Answer: dz/dt=
1
Expert's answer
2012-10-25T11:49:50-0400
x=5, y=12 then z^2=5^2+12^2 so z=+-13 z>0 so z=13
we can
write
2*z*dz=2*x*dx+2*y*dy ( divide by
dt)
2*z*dz/dt=2*x*dx/dt+2*y*dy/dt
dz/dt=( 2*x*dx/dt+2*y*dy/dt
)/2z=(2*5*2+2*12*5)/(2*13)=70/13